40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1641     Entry time: Tue Jun 14 22:44:12 2016
Author: awade 
Type: DailyProgress 
Category: Laser 
Subject: Measurement of laser output power vs current (south M126N-1064-500) 

Antonio mentioned that he thought the power at the output of the 'south' laser used to be higher.  A measurement was made 3-6 months ago of power verses laser current (and the power reported on the laser head). I think this is the post to refer to: https://nodus.ligo.caltech.edu:8081/CTN/1600 which indicates a power of 306 mW. Missing from the data is the laser current values (Antonio has these on file). The most coming directly out of the head now is ~235 mW. Even with a 10% error on the thorlabs power head (a S130S in this case), this seems like the power has depreciated.

It is good to document the state of the laser so I went back to the lab and stepped the laser through a bunch of currents and measured the power directly out of the laser head (for the south M126N-1064-500). Figure one shows the first take of data; the data has a discontinuous jump at about 1.99 A.

I thought that maybe I had made an error in taking down power or current, so went back to the lab and remeasured in finer increments.  It turns out that results were reproduced exactly, see Figure 2.  It appears that maybe we are going through a mode hop or maybe some other (maybe) temperature related jump as the laser reaches the higher powers.  Interestingly, once I tripped over the 1.99 A mark and then stepped back down the laser stuck at the slightly lower power.  The only way to 'reset' the effect was to dial the current right down, wait 5 minutes and then gradually bring it back up to the 1.99 A point.  This effect was repeatable.

 

--Also, for good measure, here are the settings as displayed on the south laser controller at the nominal 178 mW operating power

ADJ  = 0
DC = 2.08 A
DPM = 0.00V
Neon/off
LDON/off
Display 5
DT = 28.7 C
DTEC = +0.2V
LT = 34.8 C
LTEC = +0.5V
T = +34.4965
Pwr = 179 mW

Data for plots are in the attached Matlab script along with stuff to plot it.

-awade Tue Jun 14 22:37:55 2016

 

 

Attachment 1: plot20160614_MeasurmentOfSouthLaser(M126N-1064-500)PvsI.eps  19 kB  | Hide | Hide all
plot20160614_MeasurmentOfSouthLaser(M126N-1064-500)PvsI.eps
Attachment 2: plot20160614_REMEASUREOfSouthLaser(M126N-1064-500)PvsI.eps  21 kB  | Hide | Hide all
plot20160614_REMEASUREOfSouthLaser(M126N-1064-500)PvsI.eps
Attachment 3: plot20160614__MeasurmentLaserPowerVsDriverCurrent.m.zip  1 kB  Uploaded Fri Sep 15 22:03:55 2017
ELOG V3.1.3-